18 The atmospheric composition of CO2 in a locale is measured as 0.06796. Explain the possible physiological and /or health implications of residents of such an area. Use calculated values of partial pressures to justify your explanations and conclusions. Effect your calculations in a sequential manner. Your calculated values need to be accompanied with the proper units of measurements.

hi I'm lost. what calculated values is he speaking of? what measurements do I use?
Answer #1

Percent composition of CO2 = 0.067%

Atmospheric pressure = 760 mm Hg

Partial pressure = mole fraction × total pressure

Mole fraction = moles of CO2/total moles

Moles = no. Of molecules × NA (Avagadro number)

For every 100 molecules 0.067 molecules are CO2

Moles of CO2 = 0.067 × NA

Total moles = 100 × NA

Mole fraction = 0.067 NA/100NA = 67 x 10-5 = 0.00067

Partial pressure = mole fraction x atm pressure

Partial pressure = 0.00067 x 760 mmHg = 0.5092 mm Hg

Usual partial pressure of carbon dioxide = 0.3 mmHg but the calculated value is 0.5092 mmHg

The physiological or health implication of residents in that locale may suffer from headache, nausea, vomiting, visual impairment, intercostal pain, increased blood pressure and other respiratory issues etc

Slight increase in CO2 leads to disturbance in respiratory problems. Usually breathing involves in release of oxygen to the cells required and releases out the carbon dioxide.

So there is exchange of oxygen and carbon dioxide between the body and atmosphere. Thus carbon dioxide helps in feedback mechanism and regulate the rate of breathing.

High levels of CO2 leads to high levels of acid/lower pH and finally acidosis.

Increased levels of CO2 in air leads to the inability of body tissues to expel CO2 out.

Continous exposure may lead to congnitive and learning ability get disturbed in children.

Sometimes heavy exposure lead to kidney failure, bone atrophy and sometimes loss of brain function.